OpenStudy (anonymous):

Can Anyone Help Me please?!<3 Jane poured 12 cups of water into a rectangular container measuring 12 inches by 8 inches by 9 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. (3 points) Part B: How many cubic inches of water was poured into the container? (3 points) Part C: What was the height of the water in the container? Show your work. (4 points) Measurement is given as length by width by height.

OpenStudy (johnweldon1993):

Hmm...well the volume = lwh ...so what is the volume of the container? Hint *notice that the height is in centimeters...convert that to inchest first

OpenStudy (johnweldon1993):

So Volume = L*W*H V = 12 inches * 8 inches * 3.54 inches V = ?

OpenStudy (anonymous):

339.84?

OpenStudy (anonymous):

So that was all part A right.?

OpenStudy (johnweldon1993):

Right....but remember that is cubic inches... so the volume = 339.84 ...that is your answer to 1) 2) how much water did she pour in? well...12 cups of 14.44 cubic inches = 12 x 14.44 = ?

OpenStudy (anonymous):

173.28?

OpenStudy (johnweldon1993):

Mmhmm again 173.28 cubic inches though (don't forget to write units) so thats part 2)

OpenStudy (anonymous):

so part b is 173.28 cubic inches?

OpenStudy (johnweldon1993):

mmhmm that is part B

OpenStudy (anonymous):

you explain this so much better then my actual teachers ever did ;-;

OpenStudy (johnweldon1993):

For part 3) you take that number (173.28) and you divide it by the length and the width given (since you want the height) so \[\large \frac{173.28}{12 \times 8} = ?\]

OpenStudy (johnweldon1993):

Haha no problem, I'm glad it's making sense...sometimes it helps to have a second way of hearing it :)

OpenStudy (anonymous):

so thats 1.805?

OpenStudy (johnweldon1993):

That's correct! The water is 1.805 inches high

OpenStudy (anonymous):

Ughh thank you so so so much! your a life saver! your the person i need to go too from now on! lol

OpenStudy (johnweldon1993):

Haha I'll be here :) feel free to fan/message me anytime :)

OpenStudy (anonymous):

Thanks John!! (:

OpenStudy (johnweldon1993):

Of course Courtney :)